# Measurement, and Mitigation

### v1

@@ 4 Memorization at different stages
Pre-training dynamics introduce systematic biases that influence which training examples are retained. Under standard non-deterministic training regimes, characterized by data shuffling, dropout, and stochastic optimization, models tend to forget examples seen early in training unless they are revisited, as shown in Jagielski et al. (2023), Kiyomaru et al. (2024) and Leybzon and Kervadec (2024). They hypothesize that this forgetting is not simply due to limited exposure but rather the result of parameter drift, whereby updates driven by later data overwrite earlier representations. Consequently, memorization is biased toward examples encountered in the latter stages of training.
@@
This effect becomes even more pronounced as models approach the end of training. Huang et al. (2024) showed that later-stage checkpoints are more susceptible to memorizing even rare or out-of-distribution content, likely reflecting increased model capacity and representational flexibility at those stages. Extending this perspective, Biderman et al. (2023) demonstrated that memorization follows predictable scaling laws. As model size and training duration increase, specific sequences predictably transition from unmemorized to memorized. This behavior reinforces that memorization is not an artifact of overfitting but an emergent property of scale and optimization dynamics.
@@
Supervised fine-tuning influences memorization behavior, shaping both the extent and nature of information retention and exposure. Mireshghallah et al. (2022) provided a comparative analysis of fine-tuning approaches, demonstrating that head-only fine-tuning presents the highest risk of memorization, likely due to overfitting. In contrast, adapter-based fine-tuning, when constraining parameter updates (Zeng et al., 2024), reduces memorization using parameter-efficient methods, demonstrating fine-tuning can be effective for limiting privacy risks. They link differences in memorization from models fine-tuned for tasks such as summarization vs. question answering to the attention dynamics of each task, with narrower attention patterns correlating with higher memorization. From an adversarial perspective, Chen et al. (2024) demonstrated that fine-tuning can also be leveraged to extract memorized pretraining data. Using the Janus Interface, they show that targeted fine-tuning amplifies a model’s capacity to leak sensitive information, thus framing fine-tuning as a mechanism that can reactivate latent memorization vulnerabilities. Together, these studies underscore that fine-tuning critically shapes memorization patterns.
@@
Reinforcement learning as post-training. Post-training paradigms such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024), and Reinforcement Learning from Internal Feedback (RLIF) (Zhao et al., 2025) leverage reinforcement learning to train LLMs from various feedback sources. Limited research exists on memorization propagation through reinforcement learning stages. In the context of code generation, Pappu et al. (2024) found that data memorized during fine-tuning persists with high frequency in post-RLHF models, while finding minimal evidence that reward model data or reinforcement learning data becomes memorized. Critical questions remain regarding the predictability of memorization persistence between fine-tuning and reinforcement learning based on data attributes.
@@
Distillation is another common technique used in modern ML pipelines (Hinton et al., 2015; Zhao et al., 2022b), where a small ‘student’ model is trained on data produced by a larger ‘teacher’ model (typically logits). Chaudhari et al. (2025) show that bias injected adversarially into teacher models can be amplified in student models via distillation. This naturally suggests that memorization can propagate between teacher and student models, but this has not been formally analyzed.
@@
Open Questions 1. How can we reduce parameter drift to retain early training knowledge? 2. Do memorization scaling laws apply to domain-specific fine-tuned models? 3. How can we quantify and attribute memorization to pre-training vs. fine-tuning? 4. Are there identifiable factors that predict what type of data will persist as memorized between pre-training and post-training? 5. How does memorization persist (or not) from teacher to student models during distillation?
@@ 5 Detecting memorization
Divergence attack. Nasr et al. (2023) proposed a divergence attack, a prompt-based extraction method that circumvents alignment defenses in instruction-tuned LLMs, coaxing the model toward reverting to pre-aligned behavior. This divergence significantly increases the likelihood of emitting memorized training data, achieving up to 150$\times$ more verbatim sequences compared to typical user queries. Nasr et al. (2023) hypothesized that this vulnerability arises because prompts induce decoding analogous to an end-of-text token during pretraining, a context in which LLMs are known to favor high-likelihood and often memorized continuations. The attack exposes persistent memorization that may have remained latent.
@@
Prefix-based data extraction attack works by querying a model with the initial segment of a memorized sequence and observing whether it completes the rest (verbatim or approximate, depending on the specific definition used). First proposed by Carlini et al. (2021), they demonstrated this technique on GPT-2, successfully recovering hundreds of memorized examples by prompting the model with partial strings from the training set. Carlini et al. (2023) showed that longer prefixes increased the likelihood of verbatim completions by reducing ambiguity and aligning the model more strongly with memorized content. Building on Carlini et al. (2023), Li et al. (2024) introduced the LLM-PBE toolkit, a benchmark for evaluating privacy risks via prefix-based extraction. Li et al. (2024)’s findings reinforced that structured prefixes, such as email headers or document beginnings, are potent at eliciting memorized sequences.
@@
An adversarial variant of this attack was explored by Kassem et al. (2024), who used LLMs to generate candidate prefixes likely to elicit private data from a target model. Kassem et al. (2024) observed that instruction-tuned models can leak pretraining data even when the prompts diverge from the original training distribution, suggesting that memorization in LLMs may be more pervasive than what is revealed by prefix attacks. An extension of prefix attacks is soft prompting, which is explored in further detail below.
@@
Membership inference attack (MIA) (Shokri et al., 2017) has emerged as a key diagnostic tool for detecting memorization in LLMs, offering insight into whether specific data points may have been memorized during training. To circumvent the computational overhead of training shadow models, recent work has focused on black-box strategies that exploit the model’s output probabilities or perplexity for input sequences (Duan et al., 2024).
@@
A number of MIA techniques rely on loss- or likelihood-based metrics. Yeom et al. (2018) used the raw loss on a target input, assuming seen examples have lower loss. To mitigate confounding effects from input difficulty, reference-based calibration (Carlini et al., 2021) subtracts the loss from a separate reference model, aiming to isolate differences due to training exposure. Another method used in Carlini et al. (2021) is zlib entropy, which normalizes loss by the zlib-compressed size of the input sequence, using compression as a proxy for complexity. Moving beyond single-point estimates, the neighborhood attack (Mattern et al., 2023) examines local changes in the loss landscape. By generating nearby perturbations of the target input and comparing their average loss, it identifies instances where the target appears memorized. Similarly, the min-$k$prob method (Shi et al., 2023) focuses on the subset of tokens with the highest loss—under the assumption that even the most uncertain tokens in a memorized sequence will still be predicted with confidence.
@@
Yet, despite their utility in highlighting patterns of memorization, MIAs fall short when used as per-instance indicators of training data inclusion. As Zhang et al. (2025) and Duan et al. (2024) argued, MIAs lack a well-calibrated null model since one cannot feasibly train an identical model without the target input. This makes it impossible to meaningfully estimate false positive rates, undermining the statistical soundness of individual predictions. This limitation is underscored by the critique in Maini and Suri (2024), which challenges the claims made in Zhang et al. (2024). The latter introduces the PatentMIA benchmark and a divergence-based calibration method for MIAs, reporting improved reliability in detecting whether a sequence appeared in pretraining. However, Maini and Suri (2024) showed these gains are likely artifacts of distributional shifts introduced by the benchmark itself, which MIAs can exploit without truly resolving the underlying calibration problem. In light of these concerns, Zhang et al. (2025) recommended that MIAs be re-framed as tools for aggregate-level privacy auditing rather than as evidence of individual training data exposure. For stronger guarantees, they advocate for alternatives, such as data extraction attacks or canary-based MIAs, which offer more direct insights into memorization and leakage.
@@
Soft prompting techniques have emerged as powerful tools, enabling both the extraction and suppression of memorized content. By leveraging learned embeddings pre-pended to model inputs, continuous soft prompts can explicitly influence what a model reveals from its training data. For instance, Ozdayi et al. (2023) demonstrated that fixed-length continuous prompts could be trained to either amplify or diminish memorization, as measured by the extraction rates of secret sequences. Their results showed that attack prompts increased memorization leakage by up to 9.3%, while suppress prompts decreased extraction by up to 97.7%. Building on Ozdayi et al. (2023), Kim et al. (2023) introduced ProPILE, a privacy auditing framework that uses soft prompt tuning in a white-box setting to extract PII memorized by LLMs. Their learned prompts were transferable across models, suggesting consistent memorization patterns that can be systematically exploited.
@@
Dynamic soft prompting techniques extend the static nature of continuous prompts by conditioning on the input context. Wang et al. (2024) proposed prefix-conditioned prompting, where a prompt generator produces tailored soft prompts based on the input prefix. This dynamic approach allows LLMs to adapt to subtle input variations and surface context-dependent memorized completions. Wang et al. (2024) revealed that dynamic prompts significantly improved the discoverable memorization rate, surpassing both static prompt and no-prompt baselines. These results underscore the role of context-sensitive prompting in uncovering latent memorized content and highlight the broader utility of soft prompting in analyzing memorization in LLMs.
@@
Reasoning. Much of recent literature has been focused on training Large Reasoning Models (LRMs), models trained to reason before producing a final answer, as opposed to immediately producing an answer in their first token. To date, most memorization literature has been focused on quantifying memorization of specific datapoints by way of measuring whether the model can reproduce the training example. As the field moves towards reasoning models, a definition of what it means for a model to memorize a reasoning pattern may be useful for guiding reasoning research and ensuring reasoning models generalize. Huang et al. (2025) takes a step towards this by measuring whether models trained to reason on math are able to solve slightly perturbed versions of problems in their training set. The authors find that models fail to solve problems that constitute ‘hard’ perturbations, where the problem looks superficially similar to a problem in the training set but requires a different reasoning strategy. Creating benchmarks to measure memorization of such superficial reasoning patterns remains an open and useful direction for LRM research.
@@
Open Questions 1. How can we reliably detect memorization without training data access? 2. Can LLMs infer training data from context rather than direct memorization? 3. How can we detect memorization of implicit patterns and sensitive relationships? 4. How can we build benchmarks to measure whether reasoning models memorize superficial reasoning patterns or learn general reasoning principles?

### v2

@@ 4 Memorization at different stages
Pre-training dynamics introduce systematic biases that influence which training examples are retained. Under standard non-deterministic training regimes, characterized by data shuffling, dropout, and stochastic optimization, models tend to forget examples seen early in training unless they are revisited, as shown in [kiyomaru-etal-2024-comprehensive-analysis, jagielski2023measuring, leybzon2024learning]. They hypothesize that this forgetting is not simply due to limited exposure but rather the result of parameter drift, whereby updates driven by later data overwrite earlier representations. Consequently, memorization is biased toward examples encountered in the latter stages of training.
@@
This effect becomes even more pronounced as models approach the end of training. Huang, et al. [huang-etal-2024-demystifying] showed that later-stage checkpoints are more susceptible to memorizing even rare or out-of-distribution content, likely reflecting increased model capacity and representational flexibility at those stages. Extending this perspective, Biderman, et al. [biderman2023emergentpredictablememorizationlarge] demonstrated that memorization follows predictable scaling laws. As model size and training duration increase, specific sequences predictably transition from unmemorized to memorized. This behavior reinforces that memorization should not be viewed as a simple artifact of overfitting, but rather as an emergent property of scale that is heavily exacerbated by, and often inseparable from, it.
Crucially, the field’s inconsistent use of memorization metrics creates a methodological schism that threatens the validity of many empirical findings. Verbatim completion and Membership Inference Attacks (MIAs) do not merely offer different views; they can deliver contradictory verdicts on whether a model remembers or forgets. The work of Tirumala, et al.[tirumala2022memorizationoverfittinganalyzingtraining] exposes this paradox: a model can be shown to “forget” a sequence through the lens of MIAs while simultaneously reproducing it flawlessly upon prompting. The culprit is data duplication, which creates a weak, distributed membership signal that evades MIA detection while solidifying the sequence as a trivial, high-probability output. This methodological divergence is so severe that any claim about memorization dynamics, including the role of data duplication[jagielski2024] or the rate of forgetting, is effectively ill-defined without a precise specification of the metric. This necessitates a critical re-evaluation of prior results and calls for immediate community-wide efforts toward standardized memorization evaluation protocols.
@@
Supervised fine-tuning influences memorization, shaping both the extent and nature of information retention and exposure. Mireshghallah, et al. [mireshghallah-etal-2022-empirical] provided a comparative analysis of fine-tuning approaches, demonstrating that head-only fine-tuning presents the highest risk of memorization, likely due to overfitting. In contrast, adapter-based fine-tuning, when constraining parameter updates [zeng2024exploringmemorizationfinetunedlanguage], reduces memorization by using parameter-efficient methods, indicating that fine-tuning can be effective for limiting privacy risks. They link differences in memorization from models fine-tuned for tasks such as summarization vs. question answering to the attention dynamics of each task, with narrower attention patterns correlating with higher memorization. From an adversarial perspective, Nasr et al. [nasr2023scalableextractiontrainingdata] and Chen et al. [chen2024janusinterfacefinetuninglarge] demonstrated that fine-tuning can also be leveraged to extract memorized pretraining data. Using the Janus Interface, they showed that targeted fine-tuning amplifies a model’s capacity to leak sensitive information, thus framing fine-tuning as a mechanism that can reactivate latent memorization vulnerabilities. Together, these studies underscore that fine-tuning critically shapes memorization patterns.
@@
Reinforcement learning as post-training. Post-training paradigms such as RLHF [ouyang2022training], Reinforcement Learning with Verifiable Rewards (RLVR) [lambert2024t], and Reinforcement Learning from Internal Feedback (RLIF) [zhao2025learning] leverage reinforcement learning to train LLMs from various feedback sources. Limited research exists on memorization propagation through reinforcement learning stages. In the context of code generation, Pappu, et al. [pappu2024measuringmemorizationrlhfcode] found that data memorized during fine-tuning persists with high frequency in post-RLHF models, while finding minimal evidence that reward model data or reinforcement learning data becomes memorized. Critical questions remain regarding the predictability of memorization persistence between fine-tuning and reinforcement learning based on data attributes.
@@
Distillation is another common technique used in modern ML pipelines [hinton2015distilling, zhao-etal-2022-compressing], where a small ‘student’ model is trained on data produced by a larger ‘teacher’ model (typically logits). Chaudhari, et. al. [chaudhari2025cascadingadversarialbiasinjection] showed that bias injected adversarially into teacher models can be amplified in student models via distillation. This naturally suggests that memorization can propagate between teacher and student models, but this has not been formally analyzed.
Table II aims to frame memorization not as a single failure, but as a dynamic risk that transforms across the LLM lifecycle. The challenges are structured to show this evolution: the risk is first introduced during pre-training due to data integrity issues like semantic duplication; reshaped into concentrated ‘memorization circuits’ during fine-tuning; incentivized by the reward model during RLHF; and finally, propagated as a form of inheritance through distillation. This strategic breakdown clarifies that different stages present unique vulnerabilities, demanding tailored research directions, from scalable data sanitization at the start to auditable reward models at the end. This specification of research challenges intends to reveal that effective mitigation cannot be a single, post-hoc fix but requires a comprehensive defense with targeted interventions at every phase of a model’s creation.
Development stages Research Challenges Key Barriers & Research Directions
Pre-training How do mechanistic properties of training give rise to predictable memorization patterns? Barrier: The sheer scale of web data makes exhaustive analysis or perfect conceptual deduplication computationally infeasible. Current similarity metrics often fail to detect ‘invisible’ semantic duplicates, which remain a primary driver of memorization [lee2022deduplicatingtrainingdatamakes, SHSP2024].
Direction: Shift from computationally-prohibitive post-hoc attribution to proactive, data-centric risk modeling. Focus on developing lightweight, pre-training metrics to identify and flag inherently “memorizable” content. (e.g., based on information entropy, Kolmogorov complexity estimates, or structural patterns like PII)
Supervised Fine-tuning (SFT) How does fine-tuning on a small, narrow distribution of data alter the model’s pre-trained representations to either amplify latent memorization or introduce catastrophic memorization? Barrier: Parameter-Efficient Fine-Tuning (PEFT) methods like Low-Rank Adaptation (LoRA) [hu2022lora] concentrate updates into a small subset of weights, potentially creating specialized ‘memorization circuits’ that are hard to disentangle from circuits responsible for task-specific capabilities [hou2025impact].
Direction: Develop methods for provable and efficient machine unlearning tailored to the SFT context [nguyen2025survey]. Research novel regularization techniques for PEFT that penalize the verbatim memorization of SFT examples without causing catastrophic forgetting of pre-trained knowledge.
RLHF/Post-training Can the RLHF process be modeled as an adversarial game where the policy model is incentivized to find and exploit memorized knowledge from the base model to maximize reward? Barrier: The reward model acts as a black box, and its own memorized biases can implicitly encourage the policy model to reproduce undesired content [nasr2023scalableextractiontrainingdata]. The optimization process may discover that regurgitating certain memorized sequences is an effective strategy for maximizing reward [pappu2024measuringmemorizationrlhfcode].
Direction: Reframe alignment as an adversarial game to explicitly disincentivize memorization. Instead of simple preference optimization, co-develop a “regurgitation-seeking” model trained to elicit memorized content from the policy model. The primary model is then rewarded for helpfulness and resisting the adversarial agent, forcing it to learn a more robust, generalized representation of knowledge.
Distillation Under what conditions does knowledge distillation act as a regularizer that filters out stochastic, instance-specific memorization from the teacher, versus an amplifier that entrenches high-confidence memorized patterns into the student? Barrier: The student model is optimized to match the teacher’s output distribution [hinton2015distilling]. If the teacher assigns extremely high probability to a memorized sequence, the Kullback-Leibler (KL) divergence objective directly incentivizes the student to replicate this overconfidence, facilitating the transfer of memorized content [singh2025teacher, chaudhari2025cascadingadversarialbiasinjection].
Direction: Research distillation objectives that go beyond matching raw logits. This includes distilling from calibrated or temperature-scaled teacher distributions to penalize overconfidence, or augmenting the loss with a regularizing term that minimizes the student’s divergence from a ‘safe’ public dataset.
TABLE II: A Strategic Framework of Memorization Challenges Across Development Stages.
@@
Open Questions The Lifecycle of Risk: 1. How do training dynamics govern the forgetting and stabilization of memorized information? 2. How do memorization scaling laws differ between pre-training and fine-tuning? 3. How can we causally attribute a memorized output to a specific training stage? 4. What mechanisms determine whether a pre-trained memory is reinforced or forgotten during fine-tuning? 5. How can knowledge be distilled from a teacher to a student without transferring memorized data?
OQ1. The path-dependent nature of SGD dynamics means that information learned early in training can be overwritten or “forgotten” later [jagielski2023measuring, leybzon2024learning]. The core technical challenge is to model and predict this process of catastrophic forgetting as it applies to memorized sequences. A principled understanding would enable the design of ‘privacy-aware’ training curricula that strategically leverage natural forgetting to unlearn sensitive data, or conversely, to stabilize essential knowledge against parameter drift.
OQ2. A key technical question is whether the predictable neural scaling laws for memorization from pre-training [biderman2023emergentpredictablememorizationlarge] hold in the data-poor, overparameterized SFT regime where the model is a strong inductive prior. Investigating if the established log-linear relationships persist or if new, more aggressive scaling dynamics emerge is critical.
OQ3. When a model regurgitates data, attributing the memory to either the pre-training or fine-tuning stage is a fundamental challenge. The core technical barrier is the highly non-linear dependency of the fine-tuned parameters on the initial pre-trained state, which confounds standard attribution methods like influence functions [KohLiang2017Understanding]. The research frontier is to develop novel techniques, such as multi-stage causal interventions, that can trace the provenance of a specific memory by approximating the counterfactual impact of an example’s presence in either training set.
OQ4. The core technical problem is to predict whether a given memory from pre-training will persist, be reinforced, or be overwritten during fine-tuning. A leading hypothesis is that persistence is governed by the geometric alignment between the fine-tuning loss gradient and the parameter subspace encoding the memory; alignment may reinforce it, while misalignment may cause it to be overwritten or become latent [nasr2023scalableextractiontrainingdata]. Validating this requires advances in mechanistic interpretability to first identify these ‘knowledge subspaces’ and then analyze their interaction with the fine-tuning gradient field.
OQ5. The standard knowledge distillation objective, which minimizes the KL divergence between student and teacher output distributions [hinton2015distilling], poses a privacy risk. It implicitly trains the student to mimic the teacher’s overconfident, low-entropy predictions characteristic of memorized training examples [singh2025teacher]. The central research goal is to design alternative distillation objectives that can successfully transfer generalized knowledge while simultaneously filtering out these memorization artifacts.
@@ 5 Detecting memorization
Technique Key Principle Requirements & Practicality Core Limitation & Statistical Soundness
Category: Extraction-Based (Goal: Elicit Data)
Divergence Attack [nasr2023scalableextractiontrainingdata] Coaxes an aligned model to revert to its pre-aligned, base model behavior. Access: Black-box Practicality: High (cheap prompt engineering) Targets transient alignment artifacts rather than core memorization phenomena. Lacks durability and is susceptible to mitigation via model updates.
Prefix-based Data Extraction [carlini2021extractingtrainingdatalarge] Queries model with a known training data prefix to elicit verbatim completion. Access: Requires training data Practicality: High (for auditing) / Intractable (for attack) Inapplicable to black-box threat models due to its reliance on training data access. Scope is restricted to verbatim prefix completion, offering low recall for other memorization types.
Category: Classification-Based (Goal: Classify Data)
Membership Inference Attack (MIA) [shokri2017membership, fu2024membership, huang2025df, song2025mias] Classifies data as “member” or “non-member” based on model outputs (e.g., loss). Access: Black-box Practicality: Low per-instance cost, but high setup cost for calibration models. Lacks statistical validity for per-instance claims due to the intractability of constructing a true null distribution. Individual classifications are uncalibrated and unsuitable as definitive evidence.
Category: Learned-Prompting (Goal: Discover Prompt)
Soft Prompting (Amplification) [ozdayi2023controllingextractionmemorizeddata, Schulhoff2025, kim2023propile] Uses gradient descent to find an optimal continuous prompt that maximizes extraction. Access: Full White-Box Practicality: Extremely Low (prohibitively expensive) Assumes an impractical threat model. Dependency on full white-box access and prohibitive computational cost limits its applicability to research contexts, rendering it irrelevant as a practical attack vector.
TABLE III: A Comparative Analysis of Memorization Detection Techniques.
@@
[…]
Prefix-based data extraction attack works by querying a model with a strategic input and observing whether the model completes the rest of the sequence verbatim, or approximately, depending on the specific definition used. The earliest demonstration of this phenomenon was by Carlini, et al.  [carlini2021extractingtrainingdatalarge], who successfully recovered hundreds of memorized examples from GPT-2. Critically, this initial work did not have access to the model’s training set; instead, it used diverse, strategic prompts, such as the beginning-of-sequence token or internet data not believed to be in pretraining, to induce the model to diverge and generate training content. This demonstrated that memorization existed, but the prompt methodology was less direct than modern prefix attacks.
@@
The systematic methodology of prompting with the initial segment of a known training sequence was pioneered by Lee et al. [lee2022deduplicatingtrainingdatamakes] in the context of showing that deduplication reduced memorization, thereby laying the groundwork for the modern prefix attack definition. Building on this result, Carlini, et al. [carlini2023quantifyingmemorizationneurallanguage] showed that longer prefixes increased the likelihood of verbatim completions by reducing ambiguity and aligning the model more strongly with memorized content. Subsequently, Li, et al. [li2024llmpbeassessingdataprivacy] introduced the LLM-PBE toolkit, a benchmark for evaluating privacy risks via prefix-based extraction. Li, et al.’s findings [li2024llmpbeassessingdataprivacy] reinforced that structured prefixes, such as email headers or document beginnings, are potent at eliciting memorized sequences.
@@
An adversarial variant of this attack was explored by [kassem2024alpaca], who used LLMs to generate candidate prefixes likely to elicit private data from a target model. Kassem, et al. [kassem2024alpaca] observed that instruction-tuned models can leak pretraining data even when the prompts diverge from the original training distribution, suggesting that memorization in LLMs may be more pervasive than what is revealed by prefix attacks. An extension of prefix attacks is soft prompting, which is explored in further detail below.
@@
Membership inference attack (MIA) [shokri2017membership] has emerged as a key diagnostic tool for detecting memorization in LLMs, offering insight into whether specific data points may have been memorized during training. To circumvent the computational overhead of training shadow models, recent work has focused on black-box strategies that exploit the model’s output probabilities or perplexity for input sequences [duan2024membershipinferenceattackswork].
@@
A number of MIA techniques rely on loss- or likelihood-based metrics. [yeom2018privacy] used the raw loss on a target input, assuming seen examples have lower loss. To mitigate confounding effects from input difficulty, reference-based calibration [carlini2021extractingtrainingdatalarge] subtracts the loss from a separate reference model, aiming to isolate differences due to training exposure. Another method used in [carlini2021extractingtrainingdatalarge] is zlib entropy, which normalizes loss by the zlib-compressed size of the input sequence, using compression as a proxy for complexity. Moving beyond single-point estimates, the neighborhood attack [mattern2023membership] examines local changes in the loss landscape. By generating nearby perturbations of the target input and comparing their average loss, it identifies instances where the target appears memorized. Similarly, the min-$k\%$ prob method [shi2023detecting] focuses on the subset of tokens with the highest loss, under the assumption that even the most uncertain tokens in a memorized sequence will still be predicted with confidence.
@@
Yet, despite their utility in highlighting patterns of memorization, MIAs fall short when used as per-instance indicators of training data inclusion. As Duan, et al. [duan2024membershipinferenceattackswork] and Zhang, et al. [zhang2025membershipinferenceattacksprove] argued, MIAs lack a well-calibrated null model since one cannot feasibly train an identical model without the target input. This makes it impossible to meaningfully estimate false positive rates, undermining the statistical soundness of individual predictions. This limitation is underscored by the critique in [maini2024reassessingemnlp2024s], which challenges the claims made in [zhang2024pretraining]. The latter introduces the PatentMIA benchmark and a divergence-based calibration method for MIAs, reporting improved reliability in detecting whether a sequence appeared in pretraining. However, Maini, et al. [maini2024reassessingemnlp2024s] showed these gains are likely artifacts of distributional shifts introduced by the benchmark itself, which MIAs can exploit without truly resolving the underlying calibration problem. In light of these concerns, Zhang, et al. [zhang2025membershipinferenceattacksprove] recommended that MIAs be re-framed as tools for aggregate-level privacy auditing rather than as evidence of individual training data exposure. For stronger guarantees, they advocate for alternatives, such as data extraction attacks or canary-based MIAs, which offer more direct insights into memorization and leakage.
@@
Soft prompting techniques have emerged as powerful tools, enabling both the extraction and suppression of memorized content. By leveraging learned embeddings pre-pended to model inputs, continuous soft prompts can explicitly influence what a model reveals from its training data. For instance, Ozdayi, et al. [ozdayi2023controllingextractionmemorizeddata] demonstrated that fixed-length continuous prompts could be trained to either amplify or diminish memorization, as measured by the extraction rates of secret sequences. Their results showed that attack prompts increased memorization leakage by up to 9.3%, while suppress prompts decreased extraction by up to 97.7%. Building on [ozdayi2023controllingextractionmemorizeddata], Kim, et al. [kim2023propile] introduced ProPILE, a privacy auditing framework that uses soft prompt tuning in a white-box setting to extract PII memorized by LLMs. Their learned prompts were transferable across models, suggesting consistent memorization patterns that can be systematically exploited.
@@
Dynamic soft prompting techniques extend the static nature of continuous prompts by conditioning on the input context. Wang, et al. [wang-etal-2024-unlocking] proposed prefix-conditioned prompting, where a prompt generator produces tailored soft prompts based on the input prefix. This dynamic approach allows LLMs to adapt to subtle input variations and surface context-dependent memorized completions. Wang, et al. [wang-etal-2024-unlocking] revealed that dynamic prompts significantly improved the discoverable memorization rate, surpassing both static prompt and no-prompt baselines. These results underscore the role of context-sensitive prompting in uncovering latent memorized content and highlight the broader utility of soft prompting in analyzing memorization in LLMs.
Table III further summarizes the key principles, requirements and practicality, and core limitations and statistical soundness of the above detection techniques. We observe that the divergence attack and prefix-based data extraction are both extraction attacks, as their primary goal is to compel the model to generate verbatim training data as output. In stark contrast, MIAs are score-based classifiers; they do not generate data but rather classify a given input as a training member based on model outputs like loss or perplexity. Finally, soft prompting techniques represent a distinct category of learned-prompting methods, which leverages optimization to find continuous vectors that are maximally effective at triggering memorization, separating them from attacks that use discrete, human-written prompts. These three categories (extraction, classification, and learned-prompting) serve as representative pillars in the broader landscape of detection and measurement methods.
In Appendix’s Section 2.1, Table IV organizes the primary research challenges in memorization detection into three distinct areas: definitional, methodological, and practical. The framework begins with the Defining the Detection Target, which addresses the ambiguity in what constitutes memorization (e.g., verbatim vs. paraphrased). This conceptual uncertainty leads to significant hurdles in Methodological Rigor & Reliability, where the lack of standardized benchmarks and calibrated thresholds prevents consistent and comparable evaluation of detection techniques. These upstream definitional and methodological issues directly impact the practical Scope & Viability of these methods, resulting in prohibitive computational costs and low precision that make large-scale, real-world auditing infeasible.
@@
Reasoning itself is not a detection or measurement technique, but it is used to evaluate whether an LLM is genuinely reasoning or merely regurgitating memorized patterns. Much of recent literature has been focused on training Large Reasoning Models (LRMs), models trained to reason before producing a final answer, as opposed to immediately producing an answer in their first token. To date, most memorization literature has been focused on quantifying memorization of specific datapoints by way of measuring whether the model can reproduce the training example. As the field moves towards reasoning models, a definition of what it means for a model to memorize a reasoning pattern may be useful for guiding reasoning research and ensuring reasoning models generalize. Huang, et al. [huang2025mathperturbbenchmarkingllmsmath] takes a step towards this by measuring whether models trained to reason on math are able to solve slightly perturbed versions of problems in their training set. The authors find that models fail to solve problems that constitute ‘hard’ perturbations, where the problem looks superficially similar to a problem in the training set but requires a different reasoning strategy. Creating benchmarks to measure memorization of such superficial reasoning patterns remains an open and useful direction for LRM research.
@@
Open Questions 1. How can we reliably detect memorization without training data access? 2. How can we attribute a model’s output to in-context information versus parametric memory? 3. How can we detect semantic memorization? 4. How can we distinguish memorized reasoning shortcuts from generalizable problem-solving skills?
OQ1. One open challenge is building “zero-knowledge” detectors that identify memorization from intrinsic statistical artifacts (low perplexity / high confidence), without relying on differential comparisons to training data. This requires framing detection as an out-of-distribution problem, where the goal is to distinguish plausible linguistic samples from low-complexity artifacts indicative of direct replication [schwarzschild2024rethinking].
OQ2. Distinguishing parametric recall from in-context inference is a fundamental problem of causal attribution, especially in systems like RAG where the information source is ambiguous [bai2024special]. An option is to apply mechanistic interpretability tools, such as causal tracing or activation patching [meng2022locating], to experimentally isolate the contribution of the context window versus the model’s parameters.
OQ3. Current detection methods are predominantly syntactic and fail to identify the memorization of implicit semantic relationships, a key blind spot [staab2024memorizationviolatingprivacyinference]. Addressing this requires adapting knowledge probing techniques [petroni2019language] to systematically test for a suspected relational fact using a diverse set of novel, semantically equivalent queries.
OQ4. A key challenge in evaluating LLM reasoning is to determine whether a model has learned a generalizable algorithm or has merely memorized a superficial “reasoning template”[xie2024memorization]. This requires specialized benchmarks designed to test for robustness. The most promising approach is to create rigorous perturbation benchmarks that systematically vary a problem’s surface features (e.g., numbers, variable names) while preserving its underlying logical structure. A model’s failure on such perturbed examples reveals a reliance on brittle, pattern-matching behavior rather than true procedural understanding[huang2025mathperturbbenchmarkingllmsmath].